Tavinona Company issued 1,000; $1,000 bonds for a bond issue of $1,000,000. The bonds were issued for a 10-year term with a coupon rate of 4.5%. The current market rate is 3.5%. Calculate the present (market) value of one bond, then calculate the amount of the bond premium.

Multiple Choice

  • $1,473.90; $473.90

  • $1,049.09; $49.09

  • $1,090.36; $90.36

  • $1,083.76; $83.76
Answer #1

Annual interest payment = 1,000 x 4.5%

= $45

Present value of principal to be received at maturity = Par value of bond x PVF (i%, n)

= 1,000 x PVF (3.5%, 10)

= 1,000 x 0.70892

= $708.92

Present value of interest to be received periodically over the term of the bonds = Interest x PVAF (i%, n)

= 45 x PVAF (3.5%, 10)

= 45 x 8.31661

= $374.84

Present value of bond = Present value of principal to be received at maturity + Present value of interest to be received periodically over the term of the bonds

= 708.92 + 374.84

= $1083.76

Bond premium = 1083.76 - Par value of bond

= 1083.76 - 1,000

= 83.76

Fourth option is the correct option

kindly give a positive rating if you are satisfied with the solution. Please ask if you have any query related to the question, Thanks.